# Calibration of an Irradiated Prototype for the EIC Zero-Degree Calorimeter

### v1

@@ Abstract
We study the response of a prototype Zero-Degree Calorimeter (ZDC) detector to irradiation equivalent to 1011 1-MeV protons/cm2, which matches the expected exposure after one year of operation at full nominal luminosity at the future Electron-Ion Collider (EIC). The prototype consists of 563 channels and represents about 10 percent of the final ZDC design in terms of both channel count and detector volume. It was irradiated at the NASA Space Radiation Laboratory (NSRL) at Brookhaven National Laboratory (BNL) with a dose equivalent to 1011 protons/cm2. We demonstrate that, despite significant radiation damage to the SiPMs and non-uniform degradation across the detector volume, the detector can be successfully calibrated on a channel-by-channel basis using cosmic-ray data. The damage profile, similar to what is expected in the experiment, varies by an order of magnitude or more across the detector. Even for the most heavily damaged channels, the signal-to-noise ratio for a MIP signal remains above 5. This study provides a realistic test of the system’s performance under irradiation. It complements previous SiPM-specific irradiation studies and will inform the future operation of the ZDC and other detectors that use SiPM-on-tile technology.
@@ 1 Introduction
The future Electron–Ion Collider (EIC) [Accardi2016, ABDULKHALEK2022122447] to be constructed at Brookhaven National Laboratory (BNL) will be the first collider capable of delivering high-luminosity collisions of polarized electrons with polarized protons and light ions, as well as unpolarized heavy ions. The EIC will enable a broad program to study the spin and spatial structure of the nucleon and the emergent properties of nuclear matter in terms of quarks and gluons. To fully realize this program, the ePIC detector [ePIC] has been designed as a comprehensive multi-purpose apparatus with specialized subsystems in the forward and backward regions.
@@
Silicon photomultiplier (SiPM)-on-tile technology [White_2023, refId0] has been widely adopted across ePIC due to its compactness, scalability, and excellent photon detection efficiency. It is employed in both central and forward calorimetry, where the expected annual radiation levels vary by orders of magnitude, from about $10^{9}$ to $10^{12}$ 1-MeV $n_{\text{eq}}/\text{cm}^{2}$ depending on detector position [ABDULKHALEK2022122447]. Among these systems, the Zero-Degree Calorimeter (ZDC) [Milton_2025] is particularly exposed, as it detects neutral particles emitted at very small angles with respect to the beamline. The ZDC provides essential capabilities for tagging spectator neutrons in electron–ion collisions and identifying diffractive processes, and thus must maintain stable operation under sustained high radiation doses.
@@
This work continues a sequence of SiPM-on-tile calorimeter developments carried out by our group. The first-generation device, consisting of 10 layers and 40 channels, demonstrated the feasibility of the concept in a controlled beam test at Jefferson Lab [instruments7040043]. A second-generation instrument, with 20 layers and 192 channels, was deployed in the STAR experiment at RHIC, where it provided the first operational validation of the technology under real collision and radiation conditions [Zhang_2025]. The present third-generation prototype extends the program with 23 layers and 563 channels, representing approximately 10% of the final ZDC design. The prototype was exposed to controlled irradiation at the NASA Space Radiation Laboratory (NSRL) at BNL, reaching a fluence of $10^{11}$ 1-MeV protons/cm2, equivalent to one year of EIC operation at nominal luminosity. This provided a realistic, system-level irradiation test of the detector.
@@
This paper studies pedestal and minimum ionizing particle (MIP) calibrations as diagnostics of SiPM degradation and detector stability under irradiation. Pedestal shifts and broadenings directly reflect radiation-induced dark current and noise evolution, while the MIP response provides a robust anchor for gain calibration and energy reconstruction. Together, these observables provide a comprehensive picture of SiPM performance under dose. By studying their evolution before and after irradiation, we demonstrate that despite substantial and non-uniform SiPM damage, channel-by-channel calibration with cosmic-ray data remains feasible and ensures adequate signal-to-noise performance. These results provide an essential benchmark for the long-term operation of the ePIC ZDC and offer broader insight for other ePIC subsystems based on similar technology.
@@ 2 Prototype
The prototype is a sampling calorimeter with 23 active layers interleaved with 22 absorber layers, as shown in Fig. 1. Each layer consists of a 2 cm-thick steel absorber, a scintillator-tile layer, and a custom PCB with SiPMs mounted directly on it. The absorber layers are built from a $3\times 3$ array of steel blocks, while each active layer contains 25 square scintillator tiles arranged in a $5\times 5$ grid. Odd and even layers are diagonally shifted relative to each other to create smaller effective cells and improve spatial resolution, using the HEXPLIT reconstruction algorithm [Paul_2024].
@@
The steel blocks were reused from the STAR FCS upgrade [Wang_2024]. Each block is a square prism measuring $9.5\times 9.5\times 2~\mathrm{cm^{3}}$, with two pairs of blind holes on opposite faces (two per face) to accommodate dowel pins. The 2 cm thickness corresponds to $1.14$ radiation lengths ($X_{0}$) and approximately $0.5$ hadronic interaction lengths ($\lambda_{I}$).
Each active layer consists of 25 EJ-212 scintillating tiles [EJ212] wrapped in reflective ESR foil. A central dimple on the back of each tile houses a Hamamatsu S14160-1315PS 1.3-mm SiPM, air-coupled to the tile. The tiles are adhered onto a custom PCB that routes signals to adapter connectors for readout.
@@
Each PCB is equipped with two adapter connectors, each supporting up to 13 channels. Custom 1 m ribbon cables link the adapters to CAEN FERS-5200 units (model DT5202) [DT5202], with compatible ports. Each CAEN unit provides five adapter ports; nine units are used to cover 22.5 layers, with the final layer reading out only half of its channels. In total, the system reads out 507 channels.
The CAEN FERS system, based on the CITIROC-1A ASIC [citiroc], is adopted for prototyping and beam-test operation, while the final readout electronics for the EIC calorimeters, based on the EIC-custom CALOROC ASIC [Dumas:2025dhq], are still under development and not yet available. While this setup does not constitute a complete full-chain test, given the use of CITIROC-1A instead of CALOROC, it provides a first step toward validating the readout design, as similar performance is expected from the two ASICs and the signal-to-noise ratio is anticipated to be dominated by the SiPM dark current.
All CAEN units were connected, either in series or parallel, to a CAEN A3261 concentrator [DT5215] via TDLink. The concentrator controls and synchronizes all CAEN units. The concentrator itself was connected to a laptop via USB and operated with the JANUS software [Janus].
@@ 3 Analysis
Data acquisition was performed using the Janus software package to configure and control the CAEN FERS-5200 readout units. The CAEN system digitizes each input signal through two parallel amplification paths, a high-gain and a low-gain channel, enabling an extended dynamic range and allowing simultaneous sensitivity to small signals and large energy deposits. Two trigger modes were employed for this study: PTRG (periodic trigger) mode and TLOGIC (trigger logic) mode. In PTRG mode, the system generated periodic triggers independent of detector activity, enabling unbiased sampling of the electronic response. This mode was used for pedestal characterization. In TLOGIC mode, each CAEN unit was triggered independently according to configurable local logic: a trigger was issued if two or more of its connected channels recorded signals above threshold within the coincidence window. This logic was chosen to preferentially select cosmic-ray events traversing multiple tiles while suppressing random noise triggers.
@@
MIP data were obtained with cosmic rays using the in situ prototype, positioned as during the beam test. Data were collected in TLOGIC mode, with each CAEN unit operating its own local coincidence trigger. A unit would issue a trigger when two or more of its connected channels recorded signals above threshold within a predefined coincidence window. This two-hit coincidence reflected both the geometry of the instrumented region (each unit servicing at most two and a half PCBs) and the expectation that a single cosmic-ray muon would produce one hit per layer along its path.
@@ 4.1 Irradiation Profile of the Prototype
The radiation dose received by each SiPM in the prototype was inferred from post-irradiation dark current measurements. An example for a single channel, measured across several bias voltages, is shown in the left panel of Fig. 4. Using benchmark curves obtained from previous irradiation studies [huang2025], which map the relationship between SiPM dark current and accumulated proton fluence, the corresponding radiation level for this channel was determined by intersecting the measured dark current on the benchmark curve. The inferred fluence is consistent across different bias voltages, demonstrating the robustness of the method.
@@
For context, Fig. 5 shows the expected radiation levels for the ePIC detector under nominal operation [ePICFluence]. This plot corresponds to an integrated luminosity of 1 fb-1, which yields a typical fluence of approximately $10^{9}$ 1-MeV $n_{\text{eq}}/\text{cm}^{2}$ for the ZDC. Scaling linearly to 100 fb-1, corresponding to one year of running at full nominal luminosity, the annual fluence reaches $10^{11}$ 1-MeV $n_{\text{eq}}/\text{cm}^{2}$, matching the radiation dose applied to the prototype in the NSRL test. This demonstrates that the irradiation study realistically reproduces the operational radiation environment expected for the final detector, providing a solid basis for evaluating SiPM response and calibration stability.
Figure 5: Designed radiation fluence for the ePIC detector at 1 fb-1 integrated luminosity.
@@ 4.2 Pedestal Calibration
Excluding these problematic channels, the remaining channels revealed a clear layer-dependent trend. Because the channel numbering increases with layer depth, the observed linearly decreasing pedestal mean and width with channel number reflects the longitudinal radiation gradient of the calorimeter. In the front layers, the pedestal mean rises significantly, approaching 500 ADCs with widths of comparable scale, while in the rear layers both the mean and width remain stable and consistent with pre-irradiation values.

### v2

# arXiv:2512.20852v2[physics.ins-det] 28 Feb 2026

@@ Abstract
We study the response of a prototype Zero-Degree Calorimeter (ZDC) detector to irradiation equivalent to 1011 1-MeV $n_{\text{eq}}/\text{cm}^{2}$, which matches the expected exposure after one year of operation at full nominal luminosity at the future Electron-Ion Collider (EIC). The prototype consists of 563 channels and represents about 10 percent of the final ZDC design in terms of both channel count and detector volume. It was irradiated at the NASA Space Radiation Laboratory (NSRL) at Brookhaven National Laboratory (BNL) with a dose equivalent to 1011 $n_{\text{eq}}/\text{cm}^{2}$. We demonstrate that, despite significant radiation damage to the SiPMs and non-uniform degradation across the detector volume, the detector can be successfully calibrated on a channel-by-channel basis using cosmic-ray data. The damage profile, similar to what is expected in the experiment, varies by an order of magnitude or more across the detector. Even for the most heavily damaged channels, the signal-to-noise ratio for a MIP signal remains above 5. This study provides a realistic test of the system’s performance under irradiation. It complements previous SiPM-specific irradiation studies and will inform the future operation of the ZDC and other detectors that use SiPM-on-tile technology.
@@ 1 Introduction
The future Electron–Ion Collider (EIC) [2, 1] to be constructed at Brookhaven National Laboratory (BNL) will be the first collider capable of delivering high-luminosity collisions of polarized electrons with polarized protons and light ions, as well as unpolarized heavy ions. The EIC will enable a broad program to study the spin and spatial structure of the nucleon and the emergent properties of nuclear matter in terms of quarks and gluons. To fully realize this program, the ePIC detector [9] has been designed as a comprehensive multi-purpose apparatus with specialized subsystems in the forward and backward regions.
@@
Silicon photomultiplier (SiPM)-on-tile technology [17, 13] has been widely adopted across ePIC due to its compactness, scalability, and excellent photon detection efficiency. It is employed in both central and forward calorimetry, where the expected annual radiation levels vary by orders of magnitude, from about $10^{9}$ to $10^{12}$ 1-MeV $n_{\text{eq}}/\text{cm}^{2}$ depending on detector position [1]. Among these systems, the Zero-Degree Calorimeter (ZDC) [14] is particularly exposed, as it detects neutral particles emitted at very small angles with respect to the beamline. The ZDC provides essential capabilities for tagging spectator neutrons in electron–ion collisions and identifying diffractive processes, and thus must maintain stable operation under sustained high radiation doses.
@@
This work continues a sequence of SiPM-on-tile calorimeter developments carried out by our group. The first-generation device, consisting of 10 layers and 40 channels, demonstrated the feasibility of the concept in a controlled beam test at Jefferson Lab [3]. A second-generation instrument, with 20 layers and 192 channels, was deployed in the STAR experiment at RHIC, where it provided the first operational validation of the technology under real collision and radiation conditions [18]. The present third-generation prototype extends the program with 23 layers and 563 channels, representing approximately 10% of the final ZDC design. The prototype was exposed to controlled irradiation at the NASA Space Radiation Laboratory (NSRL) at BNL, reaching a fluence of $10^{11}$ 1-MeV $n_{\text{eq}}/\text{cm}^{2}$, equivalent to one year of EIC operation at nominal luminosity. This provided a realistic, system-level irradiation test of the detector.
@@
This paper studies pedestal and minimum ionizing particle (MIP) calibrations as diagnostics of SiPM degradation and detector stability under irradiation. Pedestal shifts and broadenings directly reflect radiation-induced dark current and noise evolution, while the MIP response provides a robust anchor for gain calibration and energy reconstruction. Together, these observables provide a comprehensive picture of SiPM performance as a function of radiation dose. By studying their evolution before and after irradiation, we demonstrate that despite substantial and non-uniform SiPM damage, channel-by-channel calibration with cosmic-ray data remains feasible and ensures adequate signal-to-noise performance. These results provide an essential benchmark for the long-term operation of the ePIC ZDC and offer broader insight for other ePIC subsystems based on similar technology.
@@ 2 Prototype
The prototype is a sampling calorimeter with 23 active layers interleaved with 22 absorber layers, as shown in Fig. 1. Each layer consists of a 2 cm-thick steel absorber, a scintillator-tile layer, and a custom PCB with SiPMs mounted directly on it. The absorber layers are built from a $3\times 3$ array of steel blocks, while each active layer contains 25 square scintillator tiles arranged in a $5\times 5$ grid. Odd and even layers are diagonally shifted relative to each other to create smaller effective cells and improve spatial resolution, using the HEXPLIT reconstruction algorithm [15].
@@
The steel blocks were reused from the STAR FCS upgrade [16]. Each block is a square prism measuring $9.5\times 9.5\times 2~\mathrm{cm^{3}}$, with two pairs of blind holes on opposite faces (two per face) to accommodate dowel pins. The 2 cm thickness corresponds to $1.14$ radiation lengths ($X_{0}$) and approximately $0.5$ hadronic interaction lengths ($\lambda_{I}$).
Each active layer consists of 25 EJ-212 scintillating tiles [8] wrapped in reflective ESR foil. A central dimple on the back of each tile houses a Hamamatsu S14160-1315PS 1.3-mm SiPM, air-coupled to the tile. The tiles are adhered onto a custom PCB that routes signals to adapter connectors for readout.
@@
Each PCB is equipped with two adapter connectors, each supporting up to 13 channels. Custom 1 m ribbon cables link the adapters to CAEN FERS-5200 units (model DT5202) [5], with compatible ports. Each CAEN unit provides five adapter ports; nine units are used to cover 22.5 layers, with the final layer reading out only half of its channels. In total, the system reads out 563 channels.
The CAEN FERS system, based on the CITIROC-1A ASIC [4], is adopted for prototyping and beam-test operation, while the final readout electronics for the EIC calorimeters, based on the EIC-custom CALOROC ASIC [7], are still under development and not yet available. While this setup does not constitute a complete full-chain test, given the use of CITIROC-1A instead of CALOROC, it provides a first step toward validating the readout design, as similar performance is expected from the two ASICs and the signal-to-noise ratio is anticipated to be dominated by the SiPM dark current.
All CAEN units were connected, either in series or parallel, to a CAEN A3261 concentrator [6] via TDLink. The concentrator controls and synchronizes all CAEN units. The concentrator itself was connected to a laptop via USB and operated with the JANUS software [12].
@@ 3 Analysis
Data acquisition was performed using the Janus software package to configure and control the CAEN FERS-5200 readout units. The CAEN system digitizes each input signal through two parallel amplification paths, a high-gain and a low-gain channel, enabling an extended dynamic range and allowing simultaneous sensitivity to small signals and large energy deposits. Two trigger modes were employed for this study: PTRG (periodic trigger) mode and TLOGIC (trigger logic) mode. In PTRG mode, the system generated periodic triggers independent of detector activity, enabling unbiased sampling of the electronic response [^1] The peak sensing feature of CITIROC-1A ASIC causes the measured pedestal to be slightly elevated and broadened relative to the true electronics baseline. This mode was used for pedestal characterization. In TLOGIC mode, each CAEN unit was triggered independently according to configurable local logic: a trigger was issued if two or more of its connected channels recorded signals above threshold within the coincidence window. This logic was chosen to preferentially select cosmic-ray events traversing multiple tiles while suppressing random noise triggers.
@@
MIP data were obtained with cosmic rays using the in situ prototype, positioned as during the beam test. Data were collected in TLOGIC mode, with each CAEN unit operating its own local coincidence trigger, and the SiPM bias voltage was set to 43 V both before and after irradiation. A unit would issue a trigger when two or more of its connected channels recorded signals above threshold within a predefined coincidence window of 100 ns, with the signal shaping time set to 25 ns. This two-hit coincidence reflected both the geometry of the instrumented region (each unit servicing at most two and a half PCBs) and the expectation that a single cosmic-ray muon would produce one hit per layer along its path.
@@ 4.1 Irradiation Profile of the Prototype
The radiation dose received by each SiPM in the prototype was inferred from post-irradiation dark current measurements [^2] The room temperature annealing effect is negligible based on our previous study [11]. An example for a single channel, measured across several bias voltages, is shown in the left panel of Fig. 4. Using benchmark curves obtained from previous irradiation studies [11], which map the relationship between SiPM dark current and accumulated proton fluence, the corresponding radiation level for this channel was determined by intersecting the measured dark current on the benchmark curve. The inferred fluence is consistent across different bias voltages, demonstrating the robustness of the method.
@@
For context, Fig. 5 shows the expected radiation levels for the ePIC detector under nominal operation [10]. This plot corresponds to an integrated luminosity of 1 fb-1, which yields a typical fluence of approximately $10^{9}$ 1-MeV $n_{\text{eq}}/\text{cm}^{2}$ for the ZDC. Scaling linearly to 100 fb-1, corresponding to one year of running at full nominal luminosity, the annual fluence reaches $10^{11}$ 1-MeV $n_{\text{eq}}/\text{cm}^{2}$, matching the radiation dose applied to the prototype in the NSRL test. This demonstrates that the irradiation study realistically reproduces the operational radiation environment expected for the final detector, providing a solid basis for evaluating SiPM response and calibration stability.
Figure 5: Designed radiation fluence for the ePIC detector at 1 fb-1 integrated luminosity. Figure taken from [10]
@@ 4.2 Pedestal Calibration
Excluding these problematic channels, the remaining channels revealed a clear layer-dependent trend. Because the channel numbering increases with layer depth, the observed linearly decreasing pedestal mean and width with channel number reflects the longitudinal radiation gradient of the calorimeter. In the front layers, the pedestal mean rises significantly, approaching 500 ADCs with widths of a similar magnitude, while in the rear layers both the mean and width remain stable and consistent with pre-irradiation values.